OpenStudy (anonymous):

A plumber charges $55 per hour plus a $50 service charge. Kevin hired the plumber to install a new hot water heater. What are the total charges if the plumber took 4 hours to complete the installation?

OpenStudy (anonymous):

55*4 plus 50

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

270

OpenStudy (anonymous):

hey this is easy! so this is what you do. the plumber will charge you 50 bucks no matter what, even if there's noting wrong with your system. so you know that 50 bukd is in the equation. now it says 55 bucks per hour. if he worked for 4 hrs, then he earns 55* 4. so the money he earns is (55*4) + 50

OpenStudy (anonymous):

55*4+50=220+50=270$